Which of the following production funcitons exhibits constant returns to​ scale?

A.

q​ = log(KL)

B.

q​ = K​ + L

C.

q​ =

KL0.5

D.

q​ = KL
Answer #1

Ans. B. q​ = K​ + L

Suppose both labour and capital are doubled, the
2K + 2L = 2 (K+L) = 2q
Then, output also doubles

Hence, the function exhibits CRS